OpenStudy (anonymous):

1/3x+1/4y=10 1/3x-1/2y=4 Please help me solve

OpenStudy (whpalmer4):

First thing I would do would be multiply both sides of both equations by 12 to get rid of those pesky fractions. That would give us 4x + 3y = 120 4x - 6y = 48 Now we can subtract those two equations because they have matching coefficients in the x terms: 4x + 3y = 120 - 4x - 6y = 48 ------------- 9y = 72 Which you can easily solve for y. Now you substitute the value you found for y back in one of the original equations, and solve for x.

OpenStudy (whpalmer4):

To clarify, that - 4x - 6y = 48 should be interpreted as - [4x - 6y = 48]

OpenStudy (anonymous):

Ok thank you. I have worked on this for hrs and i tried to cancel out 1/3x by multiplying by -1 i like ur way better. Getting rid of fractions was a better idea. Thanks

OpenStudy (whpalmer4):

You could also multiply 4x - 6y = 48 by -1 on both sides, then add the equations which might be a bit safer as far as sign errors go. Always check your answers in one of the original equations!

OpenStudy (whpalmer4):

If you wanted to do the substitution route (wouldn't be hard, you've already got x/3 in both equations), it might look like this: x/3 + y/4 = 10 x/3 - y/2 = 4 Rearrange one of them to give just x/3 on one side: x/3 = 4 + y/2 Now substitute that in the other equation: 4 + y/2 + y/4 = 10 2y/4 + y/4 = 6 3y = 24 etc.

OpenStudy (anonymous):

Thank you so much